Circular No. 5651

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA IN THE LARGE MAGELLANIC CLOUD 1992 William Liller, Viña del Mar, Chile, reports his discovery with Probli-

com of an apparent nova in the Large Magellanic Cloud at α = 5^{h}19^{m}.7, δ = −68^{o}57′ (equinox 1950.0), providing the following magnitudes from Tech Pan film: Nov. 4.34 UT, [15; 11.21, 10.7 (red filter). The object is visible on four additional films. P. Camilleri, Cobram, Victoria, Australia, observed the object at m_{v} = 10.2 on Nov. 12.472.

GEMINGA G. F. Bignami, P. A. Caraveo, and S. Mereghetti, Istituto di Fisica

Cosmica del CNR, Milan, communicate: “Using the SUper Seeing Imager (SUSI) at the European Southern Observatory’s New Technology Telescope, we have obtained a good seeing (0′′.7) V -filter image of the Geminga field on Nov. 4. Comparing this image to those from the Canada-France-Hawaii Telescope in 1984 January and the ESO 3.6-m reflector in 1987 January (see Bignami et al. 1987, Ap.J. 319, 358), we find evidence of proper motion for G′′, the magnitude-25.5 object suggested as the optical counterpart of Geminga also by Halpern and Tytler (1988, Ap.J. 330, 201). The apparent displacement is almost due northeast and amounts to ∼ 1′′.5 between the 1984 January and 1992 November images, or ∼ 0′′.2/yr. Confirmation comes from the 1987 January data, showing an intermediate position along the same direction. With the caution required by the faintness of the object and the difficult measurement, we submit that this is independent evidence for G′′ being close, typically 100 pc for a transverse velocity of 100 km/s, and thus for it definitely being the optical counterpart of Geminga. We note in passing that the reported motion of G′′ may also be a possible explanation for the difficulties encountered with the long-term γ-ray timing of Geminga (see IAUC 5583, 5541).”

NOVA CYGNI 1992 Visual magnitude estimates (cf. IAUC 5638): Oct. 19.13 UT, 9.8 (W.

Dillon, Missouri City, TX); 28.92, 9.5 (P. Schmeer, Bischmisheim, Germany); Nov. 5.84, 9.8 (A. Boattini, Florence, Italy); 7.87, 10.1 (S. Baroni and M. V. Zanotta, Monte Panna, Italy).

1992 November 12 Daniel W. E. Green